OpenStudy (anonymous):

Which polynomial identity will prove that 133 = 8 + 125?

OpenStudy (anonymous):

Answer choices:

OpenStudy (science0229):

notice that 8=2^3 and 125=5^3

OpenStudy (whpalmer4):

Hint: \[2^3 = 8\]\[5^3=125\]

OpenStudy (anonymous):

So it's B or D?

OpenStudy (the_fizicx99):

Well notice how 2 cubed is 8 and 5 cubed is 125

OpenStudy (anonymous):

*B or C

OpenStudy (anonymous):

So is it C?

OpenStudy (whpalmer4):

difference = subtraction sum = addition to combine 125 and 8 and get 133, which operation do you need?

OpenStudy (the_fizicx99):

Sum is when you add them together so 8 + 125 = 133 :)